OpenStudy (kay.baby.):

Will Medal and Fan!! The mass of a freight train is 9.98 ⋅ 10^6 kilograms. The mass of an aircraft carrier is 7.3 ⋅ 10^7 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

OpenStudy (andrejrusa12):

@hartnn @TheSmartOne @Jadeishere @welshfella @Nnesha @JoannaBlackwelder @ganeshie8 @JFraser @Loser66 @Vuriffy @skullpatrol @Apple @ILovePuppiesLol @imqwerty @AloneS @Ambuj @agent0smith @jhonyy9 @mathstudent55 @whpalmer4 @Wolf @Ellie201 @rayep12 @rosielopez_17 @Revircs @rbrasch @Tifferslovesyaoi @TinaSh @Ttrichards @todjunior @ttttttt @YanaSidlinskiy @yasatg @UnkleRhaukus @Unavailabilityy @userk2 @oliviad @oopsydaisy35 @poopsiedoodle @PandaNinja101 @procrastilate @pooja195 @Agl202 @Aleah54 @AAbomosalam1998 @Autumn_12001 @Ashton_Bishop @anastasiai @sweetburger @samanthagreer @something @sam_OVO @smollberru @sophiesky @Daniellelovee @Dyann.Cardona @ddoll01 @Deadpool @Finette @gudetamas @Goldendragon01 @Hayhayz @HocusPucus292333393 @hadeel @htxatlarge @jtug6 @johnsmith31613 @jjjjjjjjjjjjhujljnjb @Jaynator495 @jim_thompson5910 @jabez177 @kewlgeek555 @kidrah69 @kellyspeakslouder @k_d_1 @kitten_dude @leenathan @Lacris @lyss @ljetibo @Lyssburleson @Loli @zzr0ck3r @zepdrix @basically @breanna.rae @Batman4ever @nerdkid @noma @MarcoReus11 @mantar0078 @Makayla_Caldwell @kevin @will.h

OpenStudy (unavailabilityy):

No mass tagging, please.

OpenStudy (kay.baby.):

Can anyone help?

OpenStudy (ljetibo):

Just divide the mass of aircraft carrier/mass of freight train. and you can have your answer.

OpenStudy (kay.baby.):

it has to be in simple form

OpenStudy (ljetibo):

Well, technically it says it has to be in scientific notation it also asks how many times is the aircraft carrier heavier than the freight train which says that the aircraft carrier is the heavier one. So: $$7.3/9.3$$ is something you can get from a calculator and then $$10^7/10^6$$ is something you can get over the rules for dividing powers: $$x^7/x^5 = x^{7-5}=x^2$$ which is true if you actually write x*x*x...seven times divided by x*x*...five times. Except instead of x you write 10. Then you're still left with interpreting how much is that exactly.

OpenStudy (agent0smith):

"Approximately how many more kilograms " means you gotta use subtraction